OpenStudy (anonymous):

how do i factor 5x^2=35x? thanks

hero (hero):

5x^2 - 35x = 0 5x(x-7) = 0 5x = 0 x- 7 = 0 x = 0,7

OpenStudy (amistre64):

divide each side by 5x :)

hero (hero):

Never divide by the variable you're solving for

OpenStudy (amistre64):

lol ... im a rebel

hero (hero):

:P

hero (hero):

Dividing by 5x eliminates zero as a solution

OpenStudy (anonymous):

lol thanks